TACOS! You know who doesn't love tacos? Nobody Elliot loves tacos so much, he opened a taco stand. Like yourself, Elliot is a clever businessperson who wants to harness the power of statistics to maximize his profit. He collected some data on the products he sells. The following data represent the number of delicious goodies sold per day at Elliot's World Tacos and their prices. Price (x) Units sold (y) Item (just here for funsies, you won't use these names in your calculations) Beef taco Chicken taco Veggie taco Taco salad Chips & salsa Soda (mostly Granica) $3 $4 $6 $9 $2 $1 53 31 12 17 60 71 a. Develop a least-squares regression line and explain what the slope of the line indicates. b. Compute the coefficient of determination and comment on the strength of relationship between x and y. C. Compute the sample correlation coefficient between the price and the number of food items sold.

PLEASE SHOW DETAILED WORK ON EACH ANSWER
Answer #1

Solution: $6 $9 $2 Price cx); $3 $4 $1 Units sold(y): 53 31 12 17 60 71 (1 Develop a least squares regression line and explaiSlope = b = SSxy SSxx where, SSxx = EX-1 (EX = 147 - (25) ń (Ex)? L 9 = 147 - 625 = 147 – 104.17 9 42.83 SSxy= 42.83 SSxy = Enow, where y = LEY Inter cept (a) is given by a= y - bT 244 244 40-67 X = LEX 25 4.17 6 n a = 40.67 -(-7.417) x 4.17 40.67 tFLEX-KLASY (c) Compute the sample correlation coefficient between the price @) and number of food items Soldly) Ans Correlatiis (b) Coefficient of Determination (R2) Squared value of correlation called coefficient of determination (R2) R2= (0.9026)